Solve the first order linear inhomogeneous differential equation using the bernoulli method

xy,-2y=2x4

1
Expert's answer
2022-02-06T12:04:32-0500

"y' - 2\frac{y}{x} = 2{x^3}"

Substitution: "y = uv \Rightarrow y' = u'v + uv'"

Then

"u'v + uv' -2 \frac{{uv}}{x} = 2{x^3}"

"u'v + u\left( {v' - 2\frac{v}{x}} \right) = 2{x^3}"

Let

"v' - \frac{{2v}}{x} = 0 \Rightarrow \frac{{dv}}{{dx}} = \frac{{2v}}{x} \Rightarrow \frac{{dv}}{v} = \frac{{2dx}}{x} \Rightarrow \ln v = \ln {x^2} \Rightarrow v = {x^2}"

Then

"u'{x^2} = 2{x^3} \Rightarrow u' = 2x \Rightarrow u = {x^2} + C \Rightarrow y = uv = \left( {{x^2} + C} \right){x^2} = {x^4} + C{x^2}"

Answer: "y = {x^4} + C{x^2}"